4 1/7 x 9 5/6 in fractions
Answer by lenny460(1073)

4 1/7 x 9 5/6 in fractions

4 1/7 = 29/7
9 5/6 = 59/6
Therefore:
29/7 * 59/6
= 1711/42
= 40 31/42

Answer:
40 31/42